OpenStudy (anonymous):

Solve each quadratic equation using factoring?? Pleasee help :)?? 3x^2-x-10=0

OpenStudy (anonymous):

try \((3x-5)(x+2)=0\)

OpenStudy (anonymous):

Ok. So break down the factoring (3x + ?)(x - ?). What factors of 10 will give you a difference of -1x?

OpenStudy (theviper):

3x^2-x-10=0 3x^2-6x+5x-10=0 3x(x-2)+5(x-2)=0 (3x+5)(x-2)=0 I think @satellite73 u wrong here :\

OpenStudy (anonymous):

What are you making factoring so hard, Viper????

OpenStudy (anonymous):

he's actaaully helping me because i have to show my work for problems like these

OpenStudy (anonymous):

i didn't say it was right i said try it

OpenStudy (theviper):

oh i was mad lol

OpenStudy (anonymous):

Sorry. It just looks more complicated than it really is.

OpenStudy (anonymous):

no i actually understand it lol its okaay

OpenStudy (theviper):

:)

OpenStudy (anonymous):

look out for the next one's

OpenStudy (theviper):

ok